OpenStudy (anonymous):

To make a profit, stores "mark-up" their goods. Sears prices a $12.00 shirt with a 40% markup. How much will Sears sell the shirt?

OpenStudy (dangerousjesse):

\(12\times 1.4=?\)

OpenStudy (anonymous):

12x 0.4 I believe

OpenStudy (dangerousjesse):

No, when it's marked up, you're adding the whole and the mark up percentage together. You knock out a step by just making it \(\times 1+percent\) or \(\times 1.4\)